# Learning from Interaction: A Survey

### v1

@@ 3.3 Diversity Scaling
Scaling the diversity of data is key to building more robust and generalizable agents. Managing diversity at the task level (e.g., task difficulty, task objectives) helps agents acquire broader skills rather than overfitting to specific patterns (hu2025agentgen; huang2025r). At the environment level, exposing agents to a wide range of scenarios (e.g., different domains or tool suites) can further enhance their adaptability to novel situations. For example, representative works like AgentGen (hu2025agentgen) and AgentGym (xi2024agentgym) synthesize a wide range of heterogeneous settings that broaden the training signals agents receive. Beyond these, AgentSense (leng2025agentsense) generates diverse virtual sensor data by simulating different human personas and routines, and AgentBank (song2024agentbank) shows that training on tens of thousands of heterogeneous interaction trajectories will substantially improves generalization. Collectively, these approaches demonstrate that diversity across tasks and environments is foundational for training capable, adaptable agents.
@@ 4 Stage 2: Task Execution
In the task execution stage, after the agent takes an action, it receives an observation from the environment. Consequently, whether the agent can interact with the environment in real time (interactivity § 4.1) and whether the returned observations are consistent with real-world scenarios (realism § 4.2) are both critical to the quality of the resulting experience. Accordingly, we organize environment scaling in this stage into two directions: interactivity scaling and realism scaling, as shown next to the action space in Figure 3.
@@ 4.1 Interactivity Scaling
Despite the advent of standard protocols such as the Model Context Protocol (MCP) (anthropic2024mcp; luo2025mcp; wang2025mcp; fan2025mcptoolbench++) integrates heterogeneous data sources and tools into a unified form of context and thus greatly improves the efficiency and controllability of tool use, many datasets (fan2025mcptoolbench++; liu2024apigenautomatedpipelinegenerating; qin2023toolllm) still consist of predefined, carefully curated sequences of tool calls, even some of them include real API or MCP calls. Under this non-interactive settings, each task has a single predefined solution path. Agents are blind to intermediate tool outputs and cannot adapt subsequent tool selection based on the returned results. Consequently, agents trained on this static supervision exhibit poor generalization to novel tasks and limited diversity in solution paths (sullivan2025procedural). Recent methods (pang2025browsemaster; wang2024executable; yao2023webshopscalablerealworldweb; trivedi2024appworldcontrollableworldapps; sullivan2025procedural) start to allow agents to interactively invoke real world APIs or leverage tools via function calling or code generation, where they can adjust subsequent tool selections based on current output. Among these approaches, BrowseMaster (pang2025browsemaster) further supports parallel tool calling, expanding the typical one-tool-call-per-invocation pattern to an average of 12.11 calls per invocation, which further increases interactivity. Another promising direction uses an offline real database as the interaction environment (tongyidr; fang2025generalagentic; ye2025feedbackdriven; yao2024tau; barres2025tau2bench; prabhakar2025apigenmt; ye2025toolhop), where the agent interactively calls functions to read and write the state database. This approach strikes a practical balance between interactivity and realism and helps agents accumulate meaningful experience efficiently.
@@ 4.2 Realism Scaling
To ensure that large language model (LLM) agents can generalize effectively to complex, real-world scenarios, the training data derived from these environments should maintain real-world consistency. In tool-use environments, earlier works (qin2023toolllm; lu2025toolsandbox; sun2025zerosearch) utilize LLMs to generate the results of tool calls as simulations, avoiding monetary cost and unexpected implementation errors. However, to better ensure real-world consistency, more recent approaches start to use real APIs (song2023restgpt; wu2025webwalkerbenchmarkingllmsweb; mastouri2025makingrestapisagentready) or execute tasks in simulated environments backed by offline, real-world databases (tongyidr; fang2025generalagentic; ye2025feedbackdriven; yao2024tau; barres2025tau2bench; prabhakar2025apigenmt; ye2025toolhop). Specifically, tongyidr builds a custom tool suite and a simulated environment based on an offline Wikipedia database, where agents execute actual tool calls to directly read and write to the database, achieving lower cost and higher efficiency. Beyond tool-use scenarios, where interactions are mainly text-based, the emerging paradigm of thinking with images (su2025thinking) advocates using visual information as a dynamic, manipulable workspace for intermediate reasoning. Genie 3 (parker2025genie3) further extends realism scaling to 3D scenarios by implementing a physically grounded, real-time interactive 3D world. It preserves realistic physical properties while improving long-horizon consistency, offering a practical framework and motivating the development of more realistic environments for agents to explore.
@@
Another path to improving realism of environments is the simulation of multi-agent settings. In such contexts, agents may coordinate or compete with each other (tran2025multi; qian2024scaling; li2024smoa; zhang2024chain; kim2024llmsproducefaithfulexplanations), therefore each agent’s behavior naturally becomes part of the environment for the others. As the number of agents scales, these interactions can produce emergent social and economic phenomena such as information diffusion, opinion polarization, and herding effects (yang2025oasis; zhang2025parallelized). To reproduce such societal dynamics and improve data fidelity, multi-agent frameworks like Oasis (yang2025oasis) leverage real world social media data stored in a relational database to simulate interactive environments, enabling more realistic modeling of social processes. Meanwhile, zhang2025parallelized move beyond earlier frameworks (li2023camel; gao2024agentscope) that enforce execution order via message passing SOPs by adopting the MQTT communication protocol to support asynchronous decision making among autonomous agents, more closely simulating real-world workflows. Similarly, ARE  (andrews2025arescalingagentenvironments) decouples agent and environment clocks so that world state evolves asynchronously, and it simulates other agents’ activities by treating them as independent events in the environment’s event stream. Such environments enable the simulation of real-world societal processes and support the collection of more realistic experiential data.
@@ 5 Stage 3: Feedback
In the feedback stage, the environment assesses the trajectories collected during task execution and generates feedback signals for subsequent RL training. Scaling at this stage focuses on how feedback is provided, including its frequency and richness (density § 5.1 and granularity § 5.2), its level of automation (automation § 5.3), as well as how objectively and reliably it is delivered (objectivity § 5.4 and robustness § 5.5). Accordingly, we categorize representative scaling approaches along five dimensions: density, granularity, automation, objectivity, and robustness, as shown in Figure 4.
@@ 5.1 Density Scaling
The density of feedback refers to how frequently the environment provides evaluative feedback signals. Early environments typically provided trajectory-level rewards based on the final outcome (success or failure) (jiang2025verltool; da2025agent; qian2025toolrl; wang2025otc). Despite such sparse signals leading to more stable training, especially in tasks like mathematics or code generation (deepseekai2025deepseekr1incentivizingreasoningcapability), they are insufficient for more complex multi-step tasks. In these tasks, it is challenging to verify whether an intermediate step contributes to or hinders task completion. To address this issue and scale feedback density, more recent studies (khalifa2025processrewardmodelsthink; chae2025webshepherd; gao2024designing; ma2025steplevelrewardmodelsrewarding; park2025know) have introduced step-level process-based rewards, often combined with traditional trajectory-level outcome-based rewards, thus providing denser supervision and more detailed guidance for agent improvement. However, the mechanism behind such reward designs remains underexplored, though some interesting phenomena have been observed in existing experiments. For instance, ma2025steplevelrewardmodelsrewarding shows that step-level reward models can perform well in logical-coherent tasks such as mathematical reasoning, but they are not suitable for some natural language tasks (e.g., creative writing, policy making). gao2024designing shows that naively combining step-level rewards with trajectory-level rewards can lead to reward hacking. That is, trivial actions may get high rewards, resulting in repetitive reasoning behavior that ultimately harms the agent’s training process. They further use reward differences between adjacent steps as rewards and clip them if they exceed a predefined threshold to mitigate this problem. Apart from these, park2025know finds that even state-of-the-art reward models can be poorly calibrated, assigning overly optimistic scores to some intermediate steps. These studies collectively showcase both the potential and bottlenecks of current dense reward designs. In the future, designing reward models that are more accurately calibrated, stable in training, and interpretable will be the focus of the reward density scaling.
@@ 5.2 Granularity Scaling
Granularity scaling refers to increasing the level of detail in feedback as well as enriching the forms of feedback provided. In earlier stages, evaluative feedback typically consisted of binary signals (NIPS2017_d5e2c0ad; ibarz2018rewardlearninghumanpreferences) or a single scalar score (stiennon2022learningsummarizehumanfeedback; ouyang2022traininglanguagemodelsfollow). At finer granularities, the evaluative feedback on agents’ performances is broken down into structured components, such as a set of scores across multiple criteria, offering more informative guidance (huang2025adactrl). Building on this idea, some works use manually designed, multi-faceted rewards to improve correctness and format (qian2025toolrl; song2025r1). The Rubrics as Rewards (RaR) framework (gunjal2025rubricsrewardsreinforcementlearning; huang2025reinforcement; zhou2025breaking; zhang2025chasingtail; viswanathan2025checklist) further decomposes task requirements into tangible, human-interpretable criteria. By designing rewards as checklist-style, instance-specific rubrics, it provides a middle ground between binary correctness signals and broad preference rankings.
@@ 5.3 Automation Scaling
Automation scaling refers to the process where the feedback mechanism shifts from human supervision to automated evaluations generated by artificial intelligence. Reinforcement Learning from Human Feedback (RLHF) encompasses a wide range of frameworks, with the goal of aligning large language models (LLMs) with human preferences (Sheng_2025; hu2025openrlhfeasytousescalablehighperformance; christiano2017deep). Although this method is effective, it heavily relies on human annotators, which results in a slow, costly and difficult-to-scale feedback process. In order to achieve automated feedback and reduce the reliance on labor-intensive preference annotations, an increasing number of studies have begun to use LLMs with evaluation capabilities as automated judges to replace human evaluators, thereby forming a new paradigm called "Reinforcement Learning from AI Feedback" (RLAIF) (zhang2025generativeverifiersrewardmodeling; su2025crossingrewardbridgeexpanding; chen2025judgelrmlargereasoningmodels; lee2024rlaifvsrlhfscaling; bai2022constitutionalaiharmlessnessai). The pioneering work REWARDAGENT (peng2025agenticrewardmodelingintegrating) introduced a verification agent to simultaneously assess the factual correctness of model responses and their compliance with instructions, and combine these assessment results with basic human feedback to effectively guide model training. In contrast, ARMAP (chen2025scalingautonomousagentsautomatic) innovatively bypassed the need for a more powerful LLM to act as a judge, instead using the positive and negative trajectories generated by the LLM to train the classification model, thereby constructing an automatic and efficient reward mechanism. Rubric-based methods (gunjal2025rubricsrewardsreinforcementlearning; huang2025reinforcement), on the other hand, decompose open-ended tasks into multi-dimensional, interpretable criteria, improving the quality of rewards generated by LLMs. By integrating this automated assessment into the training environment, agents can gradually acquire higher-level capabilities without the need for continuous human supervision. However, this paradigm also magnifies risks such as the propagation of LLM’s own biases in the process of adjudication, as well as the phenomenon of rewarding hacking, which refers to the situation where intelligent agents exploit the loopholes in the reward model to strive for high scores, thereby deviating from the intended goals and even causing the model to crash. This reveals the crucial trade-off between scalability and security, and emphasizes the need for research on robust risk mitigation strategies.
@@ 5.4 Objectivity Scaling
The RLVR paradigm has achieved great success in objective and easy-to-verify domains such as mathematical reasoning and code generation (shao2024deepseekmathpushinglimitsmathematical; deepseekai2025deepseekr1incentivizingreasoningcapability). However, in many real-world scenarios such as creative writing, medical consultation, and policy making, the objectivity is subjective, open, and hard to verify. In these real environments, the feedback collected often contains many biases and noise. An environment for training agents should be able to extract accurate and high-quality feedback from such kinds of objectivity, and gradually scale from simple and verifiable tasks to difficult and hard-to-verify ones. Some recent studies have made early attempts in this direction (andrews2025arescalingagentenvironments; ma2025general; akter2025nemotron; su2025crossingrewardbridgeexpanding; yu2025rlpr; liu2025nover; gunjal2025rubricsrewardsreinforcementlearning; huang2025reinforcement). For example, Writing-Zero (jia2025writingzerobridgegapnonverifiable) uses a pairwise generative model to extract reliable and verifiable signals from subjective evaluations. Omni-Thinker (li2025omni) combines rule-based verifiable rewards with generative preference signals to form a unified multi-task RL training loop. ARE’s verifier (andrews2025arescalingagentenvironments) applies RLVR by using both hard checks (exact parameter matches) and soft checks (LLM-based semantic judgments) under strict causal order. Although these works show the feasibility of structured and verifiable rewards in open-ended environments, scaling environments to handle harder forms of objectivity still faces many challenges.

### v3

@@ 2.3 Diversity Scaling
Scaling the diversity of environments and tasks is crucial for preventing agents from overfitting to specific action patterns, thereby fostering more robust and generalizable performances (hu2025agentgen; huang2025r). Although task-level diversity scaling methods can generate tasks with different objectives and levels of complexity within a fixed environment (huang2025r), enabling the agent to gradually improve, andrews2025arescalingagentenvironments argues that simple tasks in complex environments may be more effective than complex tasks in simple ones. This suggests that greater emphasis should be placed on environment-level diversity scaling. For instance, AgentGym (xi2024agentgym) and AgentBank (song2024agentbank) utilize environments from various domains (e.g., Web, embodied AI, tool use, and programming) showing that heterogeneity improves performance on both in-domain and out-of-domain tasks. More recent advances explore to automatically construct agentic environments from either pre-defined (shi2025taskcraft; andrews2025arescalingagentenvironments; yang2025swesmithscalingdatasoftware; jain2025r2egymproceduralenvironmentshybrid) or continuously refined (hu2025agentgen; zhang2025autoenv) tool and application suites. As shown by zhang2025autoenv, although no single learning strategy can handle all cases as environment diversity scales up, the performance upper bound achievable across all environments increases consistently.
[Uncaptioned image] Future Frontiers
A key challenge for scaling diversity is the diminishing marginal utility of merely adding more heterogeneous environments. Future environment designs should therefore not only scale diversity, but develop adaptive mechanisms for selecting or composing learning strategies tailored to heterogeneous environments (zhang2025autoenv).
@@ 3 Stage 2: Task Execution
In the task execution stage, interaction takes the form of the agent taking an action and receiving an observation from the environment. We organize works towards environment scaling in this stage into two dimensions: interactivity scaling and realism scaling (Figure 3). Interactivity (§ 3.1) learns how agent performance affected by the frequency of tool interaction, while realism (§ 3.2) focuses on making observations more realistic to enhance the interaction quality.
@@ 3.1 Interactivity Scaling
Interactivity concerns whether the environment is executable and whether the agent is blind to the returns of intermediate steps. In non-interactive settings, agents are evaluated against a single static ground-truth solution path, even though there may in fact exist diverse paths to successful task completion (deng2023mind2web; kapoor2024omniact; rawles2023androidwild; qin2023toolllm). Consequently, such static supervision fails to credit alternative correct solutions and limits agents’ ability to generalize to novel tasks (sullivan2025procedural; xie2024osworld; rawles2025androidworld). Recent environment designs start to allow agents to interactively invoke real-world APIs, call functions, generate codes, or interact through GUI (wang2024executable; yao2023webshopscalablerealworldweb; trivedi2024appworldcontrollableworldapps; sullivan2025procedural) and to dynamically adapt subsequent actions based on intermediate results. Width scaling methods focus on enabling parallel tool invocation (pang2025browsemaster) and broadening the interaction dimension (e.g., facilitating flexible switching between GUI operations and tool invocations (jia2025osworldmcp)), while depth scaling methods focus on efficient tool-use context management. For example, MiroThinker (miromindteam2025mirothinker) removes weakly related tool calls from the context, thereby freeing up capacity to scale up tool-use depth and yielding consistent performance gains across various benchmarks.
A bottleneck for such scaling lies in the cost of APIs calls. To alleviate this issue, several recent advances implement an offline realistic database as the environment (tongyidr; fang2025generalagentic; ye2025feedbackdriven), using function calls to read from and write to the database state, thereby keeping the cost of scaling up interactivity within a reasonable budget while preserving some realism.
[Uncaptioned image] Future Frontiers
While scaling interactivity enables richer tool use, it also introduces problems like overlong responses, diminishing marginal returns from additional tool calls, and escalating costs (miromindteam2025mirothinker). Moreover, efficacy diminishes in scenarios requiring the combination of multiple tools (jia2025osworldmcp), especially as the emerging paradigm of thinking with images (su2025thinking) introduces multi-modal tool-call returns, making their fusion significantly more challenging. Future environment designs should therefore prioritize optimizing the efficiency and quality of heterogeneous real-world tool invocation.
@@ 3.2 Realism Scaling
Realism aims to ensure that observations obtained from the environment remain consistent with the real world. In tool-use scenarios, despite the benefits brought by the Model Context Protocol (MCP) (anthropic2024mcp; luo2025mcp; wang2025mcp; fan2025mcptoolbench++), which integrates heterogeneous tools into a unified context and greatly improves the efficiency of tool use, directly calling APIs in real environments still faces challenges of monetary cost and lack of robustness (song2023restgpt; wu2025webwalkerbenchmarkingllmsweb; mastouri2025makingrestapisagentready; prabhakar2025apigenmt). Previous model-based environment designs use LLMs as a world model (hao2023reasoninglanguagemodelplanning; ge2024worldgpt; tang2024worldcoder) or to simulate the results of tool calls (qin2023toolllm; lu2025toolsandbox; sun2025zerosearch), reducing overhead and avoiding implementation errors but inevitably sacrificing realism. Facing this dilemma, offline-database methods (tongyidr; fang2025generalagentic; ye2025feedbackdriven) make a trade-off by using a snapshot of a real-world database as the interaction environment, thereby preserving a relatively realistic simulation (e.g., tool-chain failures, asynchronous external events). In Computer-Use scenarios, earlier works mainly construct on Q&A over static webpage screenshots (mialon2023gaia; deng2023mind2web; sun2022metagui; kapoor2024omniact), rather than actually operating in real executable environments. This idealized setting makes agents brittle and prone to failure in real-world conditions (e.g. similar buttons, intrusive popup windows, and heterogeneous user interfaces) that demand context awareness and fine-grained control (xie2025scalingcomputerusegroundinguser). To mitigate these, recent advances instead adopt real executable environments that involve coding (yang2023intercodestandardizingbenchmarkinginteractive), web browser navigation (yao2023webshopscalablerealworldweb; zhou2023webarena), and application control across operating systems (xie2024osworld).
@@
When extending to the multi-agent setting, interactions are no longer limited to those between agents and tools, but also occur between agents themselves. Agents may coordinate or compete with one another, so each agent’s behavior naturally becomes part of the environment for the others (tran2025multi; qian2024scaling; li2024smoa; zhang2024chain; barres2025tau2bench). Under this setting, realism scaling concerns whether the environment provides reliable underlying infrastructure and communication mechanisms (andrews2025arescalingagentenvironments; zhang2025parallelized; li2023camel; gao2024agentscope) such that, as the number of agents grows, complex social and economic phenomena (e.g., information diffusion, opinion polarization, and herding effects (yang2025oasis; zhang2025parallelized)) can emerge and be meaningfully simulated. For instance, zhang2025parallelized adopts the MQTT communication protocol to support asynchronous decision making among autonomous agents, while andrews2025arescalingagentenvironments decouples agent and environment clocks and treats other agents’ activities as independent events, thereby allowing the world state to evolve asynchronously and supporting the collection of more realistic experiential data.
[Uncaptioned image] Future Frontiers
Realistic simulation has long been a fundamental open challenge. Most existing environments follow a ReAct-style paradigm (yao2023react), assuming that the environment remains static between an agent’s observation and action. However, information arrives as overlapping streams in real world so that perception and action should occur concurrently (andrews2025arescalingagentenvironments). Future environment designs should place greater emphasis on such asynchronous settings, while also exploring ways to unlock richer scenarios and modalities.
@@ 4 Stage 3: Feedback
In the feedback stage, the environment assesses the trajectories collected during task execution and provides feedback for subsequent training. Scaling at this stage focuses on how feedback is provided, including its frequency and richness (density § 4.1 and granularity § 4.2), its level of automation (automation § 4.3), as well as how objectively and reliably it is delivered (objectivity § 4.4 and robustness § 4.5), as shown in Figure 4.
@@ 4.1 Density Scaling
The density of feedback refers to how frequently evaluative signals are provided, typically including trajectory-level outcome-based rewards and step-level process-based rewards. Trajectory-level outcome-based rewards (jiang2025verltool; da2025agent; qian2025toolrl; wang2025otc) are relatively sparse and contribute to more stable training, especially in tasks like mathematical reasoning and code generation (deepseekai2025deepseekr1incentivizingreasoningcapability). However, it is challenging to credit whether an intermediate step contributes to or hinders task completion in more complex, multi-step tasks. Step-level process-based rewards (khalifa2025processrewardmodelsthink; chae2025webshepherd; gao2024designing; ma2025steplevelrewardmodelsrewarding; park2025know), often combined with outcome-based rewards, in contrast, provide denser supervision and more detailed guidance. Still, their benefits are not consistent, with performance in natural language tasks (e.g., creative writing) being inferior to that in logical-coherent tasks (e.g., coding), and they face the problem of reward hacking, where trivial patterns receive high rewards (ma2025steplevelrewardmodelsrewarding; gao2024designing). Efforts like PR-Clip-Delta (gao2024designing) utilize reward differences between adjacent steps as rewards, clipping them if they exceed a predefined threshold, which helps mitigate repetitive behavior and the training instability associated with step-level rewards.
[Uncaptioned image] Future Frontiers
Future reward design needs to further explore the underlying mechanisms of reward models as well as their interpretability. Efforts should aim to combine the training stability of trajectory-level rewards and the fine-grained evaluation benefits of step-level rewards, expanding their application to more domains. Meanwhile, as current reward models are mostly poorly calibrated (park2025know), assigning overly optimistic scores to some intermediate steps, future designs should focus on addressing such biases.
@@ 4.2 Granularity Scaling
Granularity scaling refers to increasing the level of detail in feedback as well as enriching the forms of feedback provided. In earlier stages, evaluative feedback typically consisted of binary signals (NIPS2017_d5e2c0ad; ibarz2018rewardlearninghumanpreferences) or a single numeric score (stiennon2022learningsummarizehumanfeedback; ouyang2022traininglanguagemodelsfollow). At finer granularities, the evaluative feedback on agents’ performances is broken down into structured components, such as a set of scores across multiple criteria like correctness and format, offering more informative guidance (huang2025adactrl; qian2025toolrl; song2025r1). The Rubrics as Rewards framework (gunjal2025rubricsrewardsreinforcementlearning; huang2025reinforcement; zhou2025breaking; zhang2025chasingtail; viswanathan2025checklist) further decomposes task requirements into tangible, human-interpretable criteria. By designing rewards as checklist-style, instance-specific rubrics, it provides a middle ground between binary correctness signals and broad preference rankings.
[Uncaptioned image] Future Frontiers
Despite rubrics at finer granularity providing more detailed information than a simple overall numeric evaluation, their predefined nature by humans and reliance on internal, maybe outdated parametric knowledge will lead agents to exploit internal patterns rather than explore external knowledge (shao2025drtulu). Future work should allow rubrics to co-evolve with the expanding boundaries of knowledge.
@@ 4.3 Automation Scaling
Automation scaling involves the shift of the feedback mechanism from slow, costly, and labor-intensive human feedback to automated evaluations. The previous Reinforcement Learning from Human Feedback (RLHF) framework heavily relies on human annotators, making it difficult to scale (Sheng_2025; hu2025openrlhfeasytousescalablehighperformance; christiano2017deep). Recent works adopt LLM-as-a-Judge, train generative reward models, or utilize agentic workflows with external tools to facilitate the feedback process. LLM-as-a-Judge (chen2025judgelrmlargereasoningmodels; zheng2023judgingllmasajudge; gunjal2025rubricsrewardsreinforcementlearning; huang2025reinforcement) directly prompts off-the-shelf LLMs to provide comparative preferences or numeric scores across multiple aspects. Training-based methods (chen2025scalingautonomousagentsautomatic; zhang2025generativeverifiersrewardmodeling) train a reward model with synthesized rationales or trajectories to classify whether user intent is fulfilled, significantly outperforming simple prompt engineering but at the cost of limited generalization and higher training costs. Agentic methods (peng2025agenticrewardmodelingintegrating) further incorporate external searching tools to help verify factual accuracy and assess instruction-following ability, offering more reliable guidance, though there is still room for improvement in workflow efficiency and reward accuracy.
[Uncaptioned image] Future Frontiers
Automation magnifies risks such as the propagation of biases inherent in LLMs, like verbosity bias (favoring longer responses), position bias (favoring answers in certain positions), egocentric bias (preferring self-generated responses), and so on (zheng2023judgingllmasajudge). It also introduces the risk of reward hacking, where intelligent agents exploit loopholes in the reward model to achieve high scores, thereby deviating from the intended goals and even causing the model to crash. Future work should focus on balancing the critical trade-off between scalability and security, with an emphasis on developing robust risk mitigation strategies as automation scales.
@@ 4.4 Objectivity Scaling
Automated verification by LLMs may inevitably introduce biases and noise, while human evaluation is also subjective and limited by constrained knowledge. Therefore, it is necessary to eliminate subjectivity in environmental feedback designs. Objectivity scaling aims to reduce these biases and make automated verifiers fairer. Despite the Reinforcement Learning with Verifiable Rewards (RLVR) paradigm (lambert2025tulu3) achieving great success in some easy-to-verify domains (e.g., mathematical reasoning, code generation (shao2024deepseekmathpushinglimitsmathematical; deepseekai2025deepseekr1incentivizingreasoningcapability)), it still falls short in creative domains (e.g., writing, medical consultation, and policy making), where ground truth is often missing, and verification remains open, making it difficult to reach a unified standard. Recent advances aim to extend RLVR to these hard-to-verify domains (andrews2025arescalingagentenvironments; ma2025general; akter2025nemotron; su2025crossingrewardbridgeexpanding; yu2025rlpr; liu2025nover; gunjal2025rubricsrewardsreinforcementlearning; huang2025reinforcement; li2025omni). For instance, BRPO (jia2025writingzerobridgegapnonverifiable) introduces a pairwise generative reward model designed to offer comparative quality scores, avoiding the direct scoring of creativity in writing materials. Works like ARE (andrews2025arescalingagentenvironments) combine parameter-matching hard checks (e.g., verifying email id) with generative rubric-based soft checks to ensure relatively fair evaluation.
[Uncaptioned image] Future Frontiers
It is struggling to provide a unified standard during evaluation in domains like creative writing. In future work, when extending to broader domains and more modalities (e.g., the aesthetic evaluation of images and videos), the challenges of subjectivity will become even more pronounced. Extending RLVR to more scenarios or finding other ways to eliminate subjectivity will remain a crucial open challenge and a key frontier in the development of more intelligent agents.